OpenStudy (anonymous):

Find the slope if possible, of the line passing through each pair of points. (4,0) and (3,8,2)

OpenStudy (chihiroasleaf):

the slope of line, \(m\), that passing two points \( \Large \left( x_1, y_1\right) \) and \( \Large \left( x_2, y_2\right) \) is \[ \Large m =\frac{y_2 - y_1}{x_2 - x_1} \] from the points that you have, what is your \( x_1, y_1, x_2, y_2 \) ?

OpenStudy (anonymous):

this is coming from my book (4,0) and (3,8,2) and I don't know what I'm doing

OpenStudy (chihiroasleaf):

so there are three numbers in the second point? is it the only question with 'three number' in one point or the whole question looks like this?

OpenStudy (anonymous):

yes, ok , but can you still help me